# Precise Constraints on the Energy Budget of WASP-121 b from its JWST NIRISS/SOSS Phase Curve

### v1

@@ Abstract
Ultra-hot Jupiters exhibit day-to-night temperature contrasts upwards of 1000 K due to competing effects of strong winds, short radiative timescales, magnetic drag, and H2 dissociation/recombination. Spectroscopic phase curves provide critical insights into these processes by mapping temperature distributions and constraining the planet’s energy budget across different pressure levels. Here, we present the first NIRISS/SOSS phase curve of an ultra-hot Jupiter, WASP-121 b. The instrument’s bandpass [0.6–2.85 m] captures an estimated 50–83% of the planet’s bolometric flux, depending on orbital phase, allowing for unprecedented constraints on the planet’s global energy budget; previous measurements with HST/WFC3 and JWST/NIRSpec/G395H captured roughly 20% of the planetary flux. Accounting for the unobserved regions of the spectrum, we estimate effective day and nightside temperatures of K and K corresponding to a Bond albedo of and a heat recirculation efficiency of . Matching the phase-dependent effective temperature with energy balance models yields a similar Bond albedo of 0.3 and a mixed layer pressure of 1 bar consistent with photospheric pressures, but unexpectedly slow winds of 0.2 km/s, indicative of inefficient heat redistribution. The shorter optical wavelengths of the NIRISS/SOSS Order 2 yield a geometric albedo of (3 upper limit of 0.175), reinforcing the unexplained trend of hot Jupiters exhibiting larger Bond than geometric albedos. We also detect near-zero phase curve offsets for wavelengths above 1.5 m, consistent with inefficient heat transport, while shorter wavelengths potentially sensitive to reflected light show eastward offsets.
@@ 1 Introduction
Ultra-hot Jupiters represent some of the most extreme natural laboratories for exploring atmospheric physics, chemistry, and dynamics, particularly the mechanisms governing energy transport. These planets are generally tidally locked due to their short orbital periods, resulting in a permanent dayside facing the host star and a nightside facing toward space, creating large day-night temperature and pressure gradients which drive vigorous atmospheric circulation (Showman & Guillot, 2002; Heng & Showman, 2015). WASP-121 b is one of the most thoroughly studied ultra-hot Jupiters (Delrez et al., 2016). This highly irradiated gas giant (, ) orbits its F6V-type host star in only 1.27 days, placing it among the largest and hottest ( K) exoplanets known (Delrez et al., 2016; Bourrier et al., 2020a). Due to its inflated size and high temperatures, WASP-121 b is an ideal target for atmospheric studies, enabling high signal-to-noise observa tions at near-infrared wavelengths.
@@
The planet’s atmosphere has been studied extensively through transit observations (Evans et al., 2018; Ben-Yami et al., 2020; Gibson et al., 2020; Sing et al., 2019; Hoeijmakers et al., 2020; Evans et al., 2016; Gapp et al., 2025) and secondary eclipse measurements (Evans et al., 2017; Mikal-Evans et al., 2019, 2020), revealing an atmosphere rich with atomic and molecular species. While transmission spectra probe the upper layers near the terminator and eclipse observations cover the full dayside, each captures only isolated portions of a complex, dynamic atmosphere. Phase curve measurements complement these techniques by revealing how heat is transported across the entire planet, enabling a more complete understanding of atmospheric dynamics.
@@
Exoplanet phase curves map global temperature distributions and atmospheric compositions across a planet’s day and night sides (Parmentier & Crossfield, 2018). Spectroscopic phase curves offer an advantage over photometric curves by probing a broader range of atmospheric pressures, providing pressure-dependent insights into advective and radiative timescales (Stevenson et al., 2014). To this end, WASP-121 b has also had a number of photometric and spectroscopic phase curve observations with instruments such as TESS (Bourrier et al., 2020a; Daylan et al., 2021), Spitzer (Morello et al., 2023), Hubble (Mikal-Evans et al., 2022), and JWST NIRSpec (Mikal-Evans et al., 2023; Evans-Soma et al., 2025). These observations have revealed a low geometric albedo and strong phase-curve modulations, marked by pronounced day-night temperature contrasts and minimal phase offsets, suggestive of a low Bond albedo and inefficient heat redistribution within the atmosphere.
@@
The efficiency of horizontal energy transport on synchronously rotating exoplanets is roughly predicted by the ratio of radiative to advective timescales, which in turn is closely tied to the planet’s irradiation temperature, (Cowan & Agol, 2011a; Komacek & Showman, 2016; Parmentier et al., 2018): radiative timescale drops steeply with increasing temperature, following a relation, so the hottest planets are expected to exhibit poor day-to-night transport. This general expectation is complicated by two additional phenomena – high-temperature molecular dissociation and magnetic drag (Bell & Cowan, 2018; Menou, 2012).
@@
At the extreme temperatures of ultra-hot Jupiter atmospheres (T >2300 K), H2 can thermally dissociate on the dayside. On the nightside, it recombines, releasing latent heat that lowers the day-night temperature gradient (Bell & Cowan, 2018). Indeed, the analysis by Mikal-Evans et al. (2022) of two full-orbit HST/WFC3 phase curves showed evidence for H2O thermal dissociation and recombination matched by theoretical predictions (Parmentier et al., 2018) and recent high-resolution studies (Pelletier et al., 2025a). However, H2O dissociation/recombination is not nearly as effective as H2 for releasing latent heat, which is 100 more potent (Bell & Cowan, 2018). Other factors, such as H- bound–free and free–free opacities shaping the continuum opacity (Arcangeli et al., 2018) and enhanced magnetic drag slowing heat transport can further complicate the energy budget (Beltz et al., 2022; Beltz & Rauscher, 2024). Magnetic drag arises due to the extreme temperatures of ultra-hot Jupiters which leads to higher levels of thermal ionization of alkali metals that couple to the planet’s magnetic field. This interaction generates Lorentz forces that resist atmospheric motion, suppressing zonal winds and slowing day-to-night heat transport (Perna et al., 2010; Menou, 2012; Tan et al., 2024; Kennedy et al., 2025).
@@
Previous phase curve observations of WASP-121 b reveal significant discrepancies between observing instruments. For instance, estimates of Bond albedo and heat recirculation efficiency vary between different instruments due to differences in inferred dayside and nightside temperatures across their respective wavelength ranges (Mikal-Evans et al., 2022; Morello et al., 2023; Dang et al., 2025). Even within similar bandpasses, conflicting results emerge—Spitzer phase curve analyses at 3.6 m and 4.5 m (Morello et al., 2023), along with a subsequent 4.5 m reanalyis (Dang et al., 2025) suggest slight westward hotspot offsets, while a more recent reanalysis at both wavelengths finds an offset consistent with zero (Davenport et al., 2025). Meanwhile, phase offsets from NIRSpec NRS1 and NRS2 indicate a slight eastward shifts (Mikal-Evans et al., 2023; Evans-Soma et al., 2025). However, uncertainties introduced by instrumental differences and correction methods make these conclusions tentative (Changeat et al., 2024).
@@
A particular strength of NIRISS/SOSS is the ability to simultaneously probe reflected light and thermal emission (Coulombe et al., 2025; Morel et al., 2025). Reflected stellar light, which can only be effectively probed at visible and near-infrared wavelengths, provides valuable insights into the presence and composition of aerosols in exoplanet atmospheres. Numerous studies of reflected light at optical wavelengths have revealed reflection on cooler, western portions of the planet where clouds are expected (Demory et al., 2013; Parmentier et al., 2016; Coulombe et al., 2025). Moreover, an albedo paradox has been noted in hot Jupiters, where low geometric albedos are observed alongside higher Bond albedos (Crossfield, 2015; Schwartz & Cowan, 2015). By capturing both reflected and emitted light, NIRISS/SOSS enables us to constrain the geometric albedo from reflected phase curves while simultaneously assessing the Bond albedo from thermal emission, thereby offering a comprehensive view of an exoplanet’s energy budget and atmospheric properties.
@@
Here, we present JWST NIRISS/SOSS observations of the phase curve of WASP-121 b, covering a spectroscopic wavelength range of 0.6–2.85 µm. This range captures most of the planet’s bolometric flux, enabling a more comprehensive energy budget analysis. Additionally, it overlaps with the wavelengths covered by TESS and Hubble and bridges the gap to the near-infrared coverage of NIRSpec/395H and Spitzer/IRAC. By focusing on the NIRISS/SOSS bandpass, our analysis provides a critical comparison to previous observations, helping to reconcile discrepancies across different datasets.
In section 2 we describe our observations and the data reduction, the light curve analysis and fitting, and energy balance models. In section 3 we present the results of all our model fits and calculations. In section 4 we disscuss the implications of our results for WASP-121 b.
@@ 2.1 Observations
We obtained a full phase curve of WASP-121 b, including two secondary eclipses and one primary transit, with JWST NIRISS/SOSS (Albert et al., 2023; Doyon et al., 2023). The observations were obtained as part of the NIRISS Exploration of the Atmospheric diversity of Transiting exoplanets (NEAT) Guaranteed Time Observation Program (GTO 1201; PI D. Lafrenière) on October 26th, 2023. Our time series observation lasted for 36.90 hours consisting of 6 groups per integration for a total of 3452 integrations. We used the SUBSTRIP256 subarray (256 × 2048 pixels) that covers the three spectral orders of the NIRISS instrument.
@@ 2.2 Reduction
We use two independent reduction pipelines to extract light curves from the raw uncalibrated images: exoTEDRF (formerly supreme-SPOON; Radica, 2024; Feinstein et al., 2023; Radica et al., 2023) and NAMELESS (Coulombe et al., 2023, 2025). The independent reductions produce white light curves that are in good agreement with each other (Figure A3). We present the exoTEDRF reduction as our fiducial reduction.
@@
For the exoTEDRF reduction, we used version 1.4.0 and followed the methodology outlined in Radica et al. (2024) and Radica et al. (2025), as well as the official exoTEDRF tutorial111https://exotedrf.readthedocs.io/en/latest/content/notebooks/tutorial_niriss-soss.html. The reduction process began with Stage 1, where we skipped the dark current subtraction step to prevent the introduction of additional noise.
@@
Spectral trace centroids were determined using the edgetrigger algorithm (Radica et al., 2022). We then extracted the spectra using a simple box aperture, leveraging exoTEDRF’s built-in optimization to determine the ideal pixel width. A width of 29 pixels was found to minimize scatter, but we opted to extract with a width of 30 pixels. A box aperture extraction was chosen because self-contamination between spectral orders was found to be negligible during commissioning (Darveau-Bernier et al., 2022; Radica et al., 2022). However, we note the absence of a GR700XD/F277W exposure, which would typically aid in identifying contaminants hidden behind the spectral trace. The F277W filter, which transmits light redder than m, has proven effective for detecting such contaminants. Without this exposure, we lack a direct means to quantify dilution effects from hidden contaminants.
2.3 Light Curve Analysis
@@ 2.3.1 Astrophysical Model
Transits and eclipses are modeled using the batman Python package (Kreidberg, 2015). At shorter wavelengths (i.e., Order 2), the data are best described by a stellar variability model in which the flux varies sinusoidally with a period of 1.13 days, consistent with the stellar rotation period (Delrez et al., 2016; Bourrier et al., 2020b). At longer wavelengths (i.e., Order 1), the variability is negligible, and we assume a constant stellar flux. We therefore model the stellar flux as
@@
where is the measured eclipse depth, describes phase variations. WASP-121 b is expected to be significantly deformed by tidal forces (Delrez et al., 2016; Daylan et al., 2021). To accurately account for the planetary flux, we incorporate the variation in the planet’s projected area as a function of orbital phase (e.g., Bell et al., 2019; Kreidberg et al., 2018). We follow Bell et al. (2019) and model our phase variations as:
where describes the intrinsic brightness variations of the planet, and accounts for changes in the planet’s projected area due to its tidal deformation. For , we adopt a second-order sinusoid model following Cowan & Agol (2008):
To model the planet’s projected area, , we assume a biaxial ellipsoid, where the polar and east-west axes are of equal length. This assumption is justified as rotational deformation is expected to be negligible compared to tidal deformation (Leconte et al., 2011a). Despite the relatively short stellar rotation period of WASP-121, the high obliquity of WASP-121 b and the dominant tidal forces make rotational deformation unlikely to contribute significantly (Delrez et al., 2016). To model the projections of the biaxial ellipsoid, we adapt equations from previous studies (Leconte et al., 2011b; Bell et al., 2019):
where is the orbital inclination, is the planetary radius along polar axis and is the planetary radius at the sub-stellar point (line connecting the planet and the star).
@@ 2.3.2 Systematics Model
The systematics model corrects for potential trends in the data caused by instrument systematics or stellar variability. We make use of the optional functionality within exoTEDRF to apply a principal component analysis (PCA) to assess the stability of the SOSS trace during the course of the observation. More details of PCA analysis for SOSS can be found in Coulombe et al. (2023). We used eigenvalues from the PCA analysis to detrend the data. In particular, we detrend PCA components 2 to 4 as the 1st component corresponds to the lightcurve containing occultations and later components have minimal impact on the data. The 2nd and 3rd components likely correspond to drifts in the position of the trace, while the 4th component likely captures the characteristic beating pattern common to all SOSS observations, driven by the thermal cycling of the telescope’s instruments.
The PCA analysis revealed the presence of a minor tilt event that occurs at the 3149th integration (see Figure A2), possibly due to a minor sudden movement of one of the telescope’s primary mirror segment. A change in flux due to the tilt is not immediately noticeable in any of the lightcurves. We nonetheless include a term in the systematics model to account for any potential consequences of the tilt event.
@@
where represents the slope of the linear trend and corresponds to the middle of the observation ((). The eigenvalues of the -th PCA component are represented by and is the amplitude used for detrending. To avoid reintroducing any long-term effects from other systematic models each PCA eigenvalue is smoothed by subtracting a rolling median with a window size of 50 (Coulombe et al., 2025). We modeled the tilt event in the light-curve fitting with a Heaviside function at the time of the tilt event () where is amplitude of the jump.

### v2

@@ Abstract
Ultra-hot Jupiters exhibit day-to-night temperature contrasts upwards of 1000 K due to competing effects of strong winds, short radiative timescales, magnetic drag, and H2 dissociation/recombination. Spectroscopic phase curves provide critical insights into these processes by mapping temperature distributions and constraining the planet’s energy budget across different pressure levels. Here, we present the first NIRISS/SOSS phase curve of an ultra-hot Jupiter, WASP-121 b. The instrument’s bandpass [0.6–2.85 m] captures an estimated 50–83% of the planet’s bolometric flux, depending on orbital phase, allowing for unprecedented constraints on the planet’s global energy budget; previous measurements with HST/WFC3 and JWST/NIRSpec/G395H captured roughly 20% of the planetary flux. Accounting for the unobserved regions of the spectrum, we estimate effective day and nightside temperatures of K and K corresponding to a Bond albedo of and a heat recirculation efficiency of . Matching the phase-dependent effective temperature with energy balance models yields a similar Bond albedo of 0.3 and a mixed layer pressure of 1 bar consistent with photospheric pressures, but unexpectedly slow winds of 0.2 km/s, indicative of inefficient heat redistribution. The shorter optical wavelengths of the NIRISS/SOSS Order 2 yield a geometric albedo of (3 upper limit of 0.175), reinforcing the unexplained trend of hot Jupiters exhibiting larger Bond albedos than geometric albedos. We also detect near-zero phase curve offsets for wavelengths above 1.5 m, consistent with inefficient heat transport, while shorter wavelengths potentially sensitive to reflected light show eastward offsets.
@@ 1 Introduction
Ultra-hot Jupiters represent some of the most extreme natural laboratories for exploring atmospheric physics, chemistry, and dynamics, particularly the mechanisms governing energy transport. These planets are generally tidally locked due to their short orbital periods, resulting in a permanent dayside facing the host star and a nightside facing toward space, creating large day-night temperature and pressure gradients which drive vigorous atmospheric circulation (Showman-Guillot2002; Heng-Showman2015). WASP-121 b is one of the most thoroughly studied ultra-hot Jupiters (Delrez2016-discovery). This highly irradiated gas giant (, ) orbits its F6V-type host star in only 1.27 days, placing it among the largest and hottest ( K) exoplanets known (Delrez2016-discovery; Bourrier2020). Due to its inflated size and high temperatures, WASP-121 b is an ideal target for atmospheric studies, enabling high signal-to-noise observations at near-infrared wavelengths.
@@
The planet’s atmosphere has been studied extensively through transit observations (Evans2018; Ben-Yami2018; Gibson2020; Sing2019; Hoeijmakers2020; Evans2016; Gapp2025-NIRSpec-transmission) and secondary eclipse measurements (Evans2017; Mikal-Evans2019; Mikal-Evans2020), revealing an atmosphere rich with atomic and molecular species. While transmission spectra probe the upper layers near the terminator and eclipse observations cover the full dayside, each captures only isolated portions of a complex, dynamic atmosphere. Phase curve measurements complement these techniques by revealing how heat is transported across the entire planet, enabling a more complete understanding of atmospheric dynamics.
@@
Exoplanet phase curves map global temperature distributions and atmospheric compositions across a planet’s day and night sides (Phasecurve-book). Spectroscopic phase curves offer an advantage over photometric curves by probing a broader range of atmospheric pressures, providing pressure-dependent insights into advective and radiative timescales (Stevenson2014). To this end, WASP-121 b has also had a number of photometric and spectroscopic phase curve observations with instruments such as TESS (Bourrier2020; Daylan2021), Spitzer (Morello2023), Hubble (Mikal-Evans2022), and JWST NIRSpec (Mikal-Evans2023; Evans-Soma2025-NIRSpec-PhaseCurve). These observations have revealed a low geometric albedo and strong phase-curve modulations, marked by pronounced day-night temperature contrasts and minimal phase offsets, suggestive of a low Bond albedo and inefficient heat redistribution within the atmosphere.
@@
The efficiency of horizontal energy transport on synchronously rotating exoplanets is roughly predicted by the ratio of radiative to advective timescales, which in turn is closely tied to the planet’s irradiation temperature, (Cowan2011-albedo; Komacek&Showman2016; Parmentier2018): radiative timescale drops steeply with increasing temperature, following a relation, so the hottest planets are expected to exhibit poor day-to-night transport. This general expectation is complicated by two additional phenomena – high-temperature molecular dissociation and magnetic drag (Bell2018; Menou2012).
@@
At the extreme temperatures of ultra-hot Jupiter atmospheres (T >2300 K), H2 can thermally dissociate on the dayside. On the nightside, it recombines, releasing latent heat that lowers the day-night temperature gradient (Bell2018). Indeed, the analysis by Mikal-Evans2022 of two full-orbit HST/WFC3 phase curves showed evidence for H2O thermal dissociation and recombination matched by theoretical predictions (Parmentier2018) and recent high-resolution studies (Pelletier2025). However, H2O dissociation/recombination is not nearly as effective as H2 for releasing latent heat, which is 100 more potent (Bell2018). Other factors, such as H- bound–free and free–free opacities shaping the continuum opacity (Arcangeli2018) and enhanced magnetic drag slowing heat transport can further complicate the energy budget (Beltz2022; Beltz2024). Magnetic drag arises due to the extreme temperatures of ultra-hot Jupiters which leads to higher levels of thermal ionization of alkali metals that couple to the planet’s magnetic field. This interaction generates Lorentz forces that resist atmospheric motion, suppressing zonal winds and slowing day-to-night heat transport (Perna2010; Menou2012; Tan2024; Kennedy2025).
@@
Previous phase curve observations of WASP-121 b reveal significant discrepancies between observing instruments. For instance, estimates of Bond albedo and heat recirculation efficiency vary between different instruments due to differences in inferred dayside and nightside temperatures across their respective wavelength ranges (Mikal-Evans2022; Morello2023; Dang2025). Even within similar bandpasses, conflicting results emerge—Spitzer phase curve analyses at 3.6 and 4.5 m (Morello2023), along with a subsequent 4.5 m reanalyis (Dang2025) suggest slight westward hot spot offsets, while a more recent reanalysis at both wavelengths finds an offset consistent with zero (Davenport2025). Meanwhile, phase offsets from NIRSpec NRS1 and NRS2 indicate a slight eastward shifts (Mikal-Evans2023; Evans-Soma2025-NIRSpec-PhaseCurve). However, uncertainties introduced by instrumental differences and correction methods make these conclusions tentative (Changeat2024).
@@
A particular strength of NIRISS/SOSS is the ability to simultaneously probe reflected light and thermal emission (Coulombe2025-LTT-NIRISS-PC; Morel2025). Reflected stellar light, which can only be effectively probed at visible and near-infrared wavelengths, provides valuable insights into the presence and composition of aerosols in exoplanet atmospheres. Numerous studies of reflected light at optical wavelengths have revealed reflection on cooler, western portions of the planet where clouds are expected (Demory2013; Parmentier2016; Coulombe2025-LTT-NIRISS-PC). Moreover, an albedo paradox has been noted in hot Jupiters, where low geometric albedos are observed alongside higher Bond albedos (Crossfield2015; Schwartz2015). By capturing both reflected and emitted light, NIRISS/SOSS enables us to constrain the geometric albedo from reflected phase curves while simultaneously assessing the Bond albedo from thermal emission, thereby offering a comprehensive view of an exoplanet’s energy budget and atmospheric properties.
@@
Here, we present JWST NIRISS/SOSS observations of the phase curve of WASP-121 b, covering a spectroscopic wavelength range of 0.6–2.85 µm. This range captures most of the planet’s bolometric flux, enabling a more comprehensive energy budget analysis. Additionally, it overlaps with the wavelengths covered by TESS and Hubble and bridges the gap to the near-infrared coverage of NIRSpec/G395H and Spitzer/IRAC. By focusing on the NIRISS/SOSS bandpass, our analysis provides a critical comparison to previous observations, helping to reconcile discrepancies across different datasets.
@@ 2.1 Observations
We obtained a full phase curve of WASP-121 b, including two secondary eclipses and one primary transit, with JWST NIRISS/SOSS (Albert2023; Doyon2023). The observations were obtained as part of the NIRISS Exploration of the Atmospheric diversity of Transiting exoplanets (NEAT) Guaranteed Time Observation Program (GTO 1201; PI D. Lafrenière) on October 26th, 2023. Our time series observation lasted for 36.90 hours consisting of 6 groups per integration for a total of 3452 integrations. We used the SUBSTRIP256 subarray (256 × 2048 pixels) that covers the three spectral orders of the NIRISS instrument.
@@ 2.2 Reduction
We use two independent reduction pipelines to extract light-curves from the raw uncalibrated images: exoTEDRF (formerly supreme-SPOON; exotedrf; Feinstein2023-WASP39b; Radica2023-AwesomeSOSS) and NAMELESS (Coulombe2023; Coulombe2025-LTT-NIRISS-PC). The independent reductions produce white light-curves that are in good agreement with each other (Figure C3). We present the exoTEDRF reduction as our fiducial reduction.
@@
For the exoTEDRF reduction, we used version 1.4.0 and followed the methodology outlined in radica_2024_muted and radica_2025_promise, as well as the official exoTEDRF tutorial111https://exotedrf.readthedocs.io/en/latest/content/notebooks/tutorial_niriss-soss.html. The reduction process began with Stage 1, where we skipped the dark current subtraction step to prevent the introduction of additional noise.
@@
Spectral trace centroids were determined using the edgetrigger algorithm (Radica2022). We then extracted the spectra using a simple box aperture, leveraging exoTEDRF’s built-in optimization to determine the ideal pixel width. A width of 29 pixels was found to minimize scatter, but we opted to extract with a width of 30 pixels. A box aperture extraction was chosen because self-contamination between spectral orders was found to be negligible during commissioning (Darveau-Bernier2022; Radica2022). However, we note the absence of a GR700XD/F277W exposure, which would typically aid in identifying contaminants hidden behind the spectral trace. The F277W filter, which transmits light redder than m, has proven effective for detecting such contaminants. Without this exposure, we lack a direct means to quantify dilution effects from hidden contaminants.
2.3 Light-Curve Analysis
@@ 2.3.1 Astrophysical Model
Transits and eclipses are modeled using the batman Python package (batman). At shorter wavelengths (i.e., Order 2), the data are best described by a stellar variability model in which the flux varies sinusoidally with a period of 1.13 days, consistent with the stellar rotation period (Delrez2016-discovery; Bourrier2020-stellar-rot). At longer wavelengths (i.e., Order 1), the variability is negligible, and we assume a constant stellar flux. We therefore model the stellar flux as
@@
where is the measured eclipse depth and describes phase variations. WASP-121 b is expected to be significantly deformed by tidal forces (Delrez2016-discovery; Daylan2021). To accurately account for the planetary flux, we incorporate the variation in the planet’s projected area as a function of orbital phase (e.g., Bell2019; Kreidberg2018). We follow Bell2019 and model our phase variations as:
where describes the intrinsic brightness variations of the planet, and accounts for changes in the planet’s projected area due to its tidal deformation. For , we adopt a second-order sinusoid model following Cowan&Agol2008:
To model the planet’s projected area, , we assume a biaxial ellipsoid, where the polar and east–west axes are of equal length. This assumption is justified as rotational deformation is expected to be negligible compared to tidal deformation (Leconte2011a). Despite the relatively short stellar rotation period of WASP-121, the high obliquity of WASP-121 b and the dominant tidal forces make rotational deformation unlikely to contribute significantly (Delrez2016-discovery). To model the projections of the biaxial ellipsoid, we adapt equations from previous studies (Leconte2011b; Bell2019):
where is the orbital inclination, is the planetary radius along polar axis and is the planetary radius at the substellar point (line connecting the planet and the star).
@@ 2.3.2 Systematics Model
The systematics model corrects for potential trends in the data caused by instrument systematics or stellar variability. We make use of the optional functionality within exoTEDRF to apply a principal component analysis (PCA) to assess the stability of the SOSS trace during the course of the observation. More details of PCA analysis for SOSS can be found in Coulombe2023. We used eigenvalues from the PCA analysis to detrend the data. In particular, we detrend PCA components 2–4 as the first component corresponds to the light-curve containing occultations and later components have minimal impact on the data. The second and third components likely correspond to drifts in the position of the trace, while the fourth component likely captures the characteristic beating pattern common to all SOSS observations, driven by the thermal cycling of the telescope’s instruments.
The PCA analysis revealed the presence of a minor tilt event that occurs at the 3149th integration (see Figure C2), possibly due to a minor sudden movement of one of the telescope’s primary mirror segments. A change in flux due to the tilt is not immediately noticeable in any of the light-curves. We nonetheless include a term in the systematics model to account for any potential consequences of the tilt event.
@@
where represents the slope of the linear trend and corresponds to the middle of the observation ((). The eigenvalues of the -th PCA component are represented by and is the amplitude used for detrending. To avoid reintroducing any long-term effects from other systematic models each PCA eigenvalue is smoothed by subtracting a rolling median with a window size of 50 (Coulombe2025-LTT-NIRISS-PC). We modeled the tilt event in the light-curve fitting with a Heaviside function at the time of the tilt event () where is amplitude of the jump.